OpenStudy (anonymous):

A dog at point A goes in pursuit of a fox 30 metres away. The dog takes a leap of 2 m against 1m long leaps of the fox. If the dog makes two leaps to the fox’s three, at what distance from A will the dog catch up with the fox?

OpenStudy (anonymous):

30/2=15

OpenStudy (anonymous):

2-1=1

OpenStudy (anonymous):

2x=3 so x=3/2

OpenStudy (anonymous):

afox has already covered 30 meters so 30/1 =30 for fox and 30/2=15

OpenStudy (anonymous):

Each leap of the dog is 2 m. Each leap of the fox is 1 m. If the dog makes two leaps to the fox’s three. This means that each time the dog advances 4 m, the fox advances 3 m.

OpenStudy (anonymous):

so distance covered by dog=1/60

OpenStudy (anonymous):

Each 4 m of the dog movement, the dog is 1 m closer. The dog has to get 30 m closer. How many 4 m advances must the dog make to be 30 m closer?

OpenStudy (anonymous):

30/4=15/2

OpenStudy (anonymous):

in 15 leaps the dog would the distance 30meters

OpenStudy (anonymous):

When the dog advances 4 m, it's 1 m closer. When the dog advances 2*4 m, it's 2*1 m closer. When the dog advances 3*4 m, it's 3*1 m closer. When the dog advances 4*4 m, it's 4*1 m closer. etc. When the dog advances ___*4 m, it's 30*1 m closer. What goes in the blank?

OpenStudy (anonymous):

120m

OpenStudy (anonymous):

Correct. Since each 4 m of the dog means catching up 1 m, it takes 30 * 4m of the dog advancing to catch up all 30 m. 30 * 4m = 120 m. That is it.

OpenStudy (anonymous):

hmmmmmmmmmm..

OpenStudy (anonymous):

Did you post a problem about two runners running around a circular track?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh, I see that it was answered already.